OpenStudy (anonymous):

are the set of functions sin^2(x), cos^2(x), 4 linearly independent?

OpenStudy (anonymous):

do you know what linear independence means?

OpenStudy (anonymous):

graph them and then make your answer from the graph

OpenStudy (anonymous):

\[c_1\sin^2x + c_2\cos^2x + c_34 = 0\] for any c != 0

OpenStudy (anonymous):

@mastermindkakashi what should i look for in the graph?

OpenStudy (anonymous):

see if they touch or not

OpenStudy (freckles):

@iamheretohelp c1,c1,c3 are real numbers right?

OpenStudy (freckles):

think about the identity sin^2(x)+cos^2(x)=1

OpenStudy (freckles):

think about setting c1 and c2 so you can use that identity